# Another look at inference after prediction

### v5

@@ 3 Efficiency analysis
For the remainder of this paper, we study a modified version of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
@@
The only difference between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that all of the data, rather than the unlabeled data, is used to estimate $\bm{\gamma}^{*}$ in the augmentation term (i.e., $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ is replaced by $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$). When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$. In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix, specifically Section B.3, for a detailed analysis). Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our asymptotic analysis. We provide general results for Z-estimation and use linear regression as a running example throughout our discussion to further cultivate intuition.
@@ 3.1 Asymptotic analysis of the PPI estimator
To study the asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we present its influence function expansion in Theorem 1. The detailed derivations are deferred to the Appendix and follow those provided for the original PPI estimator.
@@
where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=\mathbb{E}\left[\frac{\partial\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})}{\partial\bm{\beta}}\right]^{-1}$. Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
We use Theorem 1 to compare the asymptotic variance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. The result in Corollary 1.1 shows that the improvement of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ increases with the proportion of unlabeled data provided that $\widehat{Y}$ is of sufficient quality.
@@
Let ${\bf{C}}_{11}=\mathbb{E}\left[\bm{\phi}(Y,X;\bm{\beta}^{*})\bm{\phi}(Y,X;\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}\right]$, ${\bf{C}}_{12}=\mathbb{E}\left[\bm{\phi}(Y,X;\bm{\beta}^{*})\bm{\phi}(\widehat{Y},X;\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}\right]$, ${\bf{C}}_{22}=\mathbb{E}\left[\bm{\phi}(\widehat{Y},X;\bm{\gamma}^{*})\bm{\phi}(\widehat{Y},X;\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]$, and ${\bf{B}}=\mathbb{E}\left[\frac{\partial\bm{\phi}(\widehat{Y},X;\bm{\gamma}^{*})}{\partial\bm{\gamma}}\right]^{-1}$. The difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right)\left(2{\bf{A}}{\bf{C}}_{12}-{\bf{B}}{\bf{C}}_{22}\right){\bf{B}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2{\bf{A}}{\bf{C}}_{12}\succ{\bf{B}}{\bf{C}}_{22}$.
We further break down this result in the case of a linear regression model.
Example 1 (Linear Regression).
Let $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$, $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, and $\bm{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$. Following Corollary 1.1, the difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
@@
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. Moreover, a sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
@@
This condition coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on $\bf{X}$}. This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al 21 and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix, Section A.4). This insight can be better understood in the case of homoskedasticity.
Specifically, suppose that $\mathbb{E}(\epsilon_{Y}^{2}\mid{\bf{X}})=\sigma_{Y}^{2}$, $\mathbb{E}(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}})=\sigma_{\widehat{Y}}^{2}$, and $\mathbb{E}(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}})=\sigma_{Y,\widehat{Y}}$,
@@
where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the partial correlation of the residuals for $Y$ and $\widehat{Y}$ given $\bf{X}$}. Thus, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for $\bf{X}$}, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function. The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.
@@ 3.2 Asymptotic analysis of the CC estimator
To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we draw on the landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators.27 While connections between the PPI estimator and AIPW have been made in the context of mean estimation21, we explictly leverage this framework to introduce a simple modification of the original PPI estimator with theoretical guarantees of improved efficiency. In the current setting, where the goal is to estimate $\bm{\beta}^{*}$ with $Y$ partially observed and $\widehat{Y}$ fully observed, AIPW estimators have influence functions of the form
${\bf{IF}}(Y,\widehat{Y},{\bf{X}};\bm{\omega})={\bf{A}}\left[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\frac{R}{\pi}+\bm{\omega}(\widehat{Y},{\bf{X}})\frac{\pi-R}{\pi}\right],$
with $\bm{\omega}(\widehat{Y},{\bf{X}})$ being a function of $\widehat{Y}$ and $\bf{X}$}. The optimal choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ with respect to asymptotic variance27 is
@@
By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$. This is a good approximation, for example, when $\bf{Z}$} contains $\bf{X}$}, the ML model captures the true regression function, and the focus is in fitting a generalized linear model with the canonical link function, $g(\cdot)$. In this case, $\bm{\phi}(Y,X;\bm{\beta})={\bf{X}}[Y-g(\bm{\beta}^{{\sf\scriptscriptstyle{T}}}{\bf{X}})]$ and $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$. However, this is also a scenario where naive inference will yield consistent estimation that is more efficient than inference based on the PPI estimator. More broadly, the PPI estimator is not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the quality of the prediction is insufficient. A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not lose efficiency from incorporating unlabeled data into estimation.
To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis. However, it is often difficult to specify an appropriate model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown. Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed. To overcome these difficulties, Chen and Chen 25 proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$. Their approach is motivated by the double-sampling design which, apart from differences in terminology (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here. The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
@@
Relating this estimator back to the PPI estimator, the high-level idea is analogous, but $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is approximated by the linear projection of $\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})$ on $\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$ rather than $\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$ itself, which effectively protects against potential inadequacies in $\widehat{Y}$ that can lead to efficiency loss.
The corresponding estimator, which we call the CC estimator, can be written as
@@
${{\bf{W}}}^{\text{\tiny{CC}}}={\bf{A}}{\text{Cov}}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{{\text{Var}}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}{\bf{B}}^{-1}$
obtained with $\mathscr{L}$ and $\mathscr{U}$. Therefore, a simple way to ensure that an estimator with the same structure as the PPI estimator has asymptotic variance no greater than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term. We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$.
@@
Following Chen and Chen 25,
@@
The subsequent corollaries apply Theorem 2 to show our two key results:(1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is optimal with respect to efficiency in the class of estimators with a weighted linear augmentation term. These results are formally stated in Corollaries 2.1 and 2.2 and show that the CC method is preferable to the PPI method.
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}{\bf{C}}_{12}{\bf{C}}_{22}^{-1}{\bf{C}}_{12}{\bf{B}}\succeq{\bf{0}}.$

### v6

@@ 3 Efficiency analysis
As PB inference methods suppose the availability of at least a small labeled data set, a natural baseline for comparison is the estimator based on $\mathscr{L}$ only, $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. For a PB inference method to be preferred over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it should not only yield valid inference, but also achieve gains in statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the prediction is informative of $Y$ and perform no worse when $\widehat{Y}$ is uninformative. For the remainder of this paper, we study the latter property and analyze a modified version of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
@@
The only difference between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that all of the data, rather than the unlabeled data, is used to estimate $\bm{\gamma}^{*}$ in the augmentation term (i.e., $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ is replaced by $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$). When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$. In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix, specifically Section B.3, for a detailed analysis). Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our asymptotic analysis.
@@ 3.1 Asymptotic analysis of the PPI estimator
To study the asymptotic propeties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we first derive its influence function expansion in Theorem 1. The detailed derivations are deferred to the Appendix.
@@
where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}$. Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
We use Theorem 1 to compare the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Corollary 1.1 shows that the benefit of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ scales directly with the predictiveness of $\widehat{Y}$ for $Y$ and inversely with the proportion of labeled data $\pi$.
@@
Let $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$, $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, and $\bm{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$. The difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
@@
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
@@
The sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on $\bf{X}$}. This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al angelopoulos2023prediction and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix, Section A.4). Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
Example 1 (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ under homoskedasticity).
In the homoskedastic setting where $\mathbb{E}(\epsilon_{Y}^{2}\mid{\bf{X}})=\sigma_{Y}^{2}$, $\mathbb{E}(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}})=\sigma_{\widehat{Y}}^{2}$, and $\mathbb{E}(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}})=\sigma_{Y,\widehat{Y}}$,
@@
where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of the residuals for $Y$ and $\widehat{Y}$ given $\bf{X}$}. Thus $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for $\bf{X}$}, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function. The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.
@@ 3.2 Asymptotic analysis of the CC estimator
To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we draw on the landmark paper that introduced the class of augmented inverse probability weighted (AIPW) estimators. robins1994estimation In the current setting, where the goal is to estimate $\bm{\beta}^{*}$ with $Y$ partially observed and $\widehat{Y}$ fully observed, these estimators have influence functions of the form
$\text{IF}(Y,\widehat{Y},{\bf{X}};\bm{\omega})={\bf{A}}\left[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\frac{R}{\pi}+\bm{\omega}(\widehat{Y},{\bf{X}})\frac{\pi-R}{\pi}\right],$
with $\bm{\omega}(\widehat{Y},{\bf{X}})$ being a function of $\widehat{Y}$ and $\bf{X}$}. The optimal choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ with respect to asymptotic variance is
@@
By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is thus an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$. The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}$. A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not be negatively impacted by incorporating unlabeled data into estimation. This raises concerns about the practical utility of the PPI estimator.
To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis. However, it is often difficult to specify an appropriate model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown. Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed. To overcome these difficulties, Chen and Chen chen2000unified proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$. Their approach is motivated by the double-sampling design which, apart from differences in terminology (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here. The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
@@
The corresponding estimator can be written as
@@
${{\bf{W}}}^{\text{\tiny{CC}}}={\bf{A}}{\text{Cov}}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{{\text{Var}}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}{\bf{A}}^{-1}$
obtained with $\mathscr{L}$. Therefore, a simple way to ensure that an estimator with the same structure as the PPI estimator has asymptotic variance no greater than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term. We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$.
@@
Following Chen and Chen chen2000unified,
@@
The subsequent corollaries apply Theorem 2 to show our two key results: (1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ has the highest asymptotic efficiency within the class of estimators with a weighted linear augmentation term. These results suggest that the CC method is preferable to the PPI method for gaining efficiency.
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$